OpenStudy (anonymous):

A circle has a diameterof 8 feet. what is the approximate area? A)12.56ft(squared) C)50.24ft(squared) B)25.12ft(squared) D)200.96ft(squared)

OpenStudy (anonymous):

C)50.24ft(squared)

OpenStudy (anonymous):

radius is 4 4^2 = 16 16 x 3 = 48

OpenStudy (anonymous):

The equation for the area of a circle is \[\pi r^2\] and the radius of the circle is the diameter/2. so you have: \[\pi 4^2 = 50.24ft^2\]

OpenStudy (anonymous):

\[\pi = 3.14159...\]